# Geometry of polynomial neural networks

### v1

@@ 6.2. Gradient-based optimization
6.3. Connecting the learning degree to empirical observations.
We consider a polynomial neural network with architecture $\mathbf{d}=(2,2,3)$ and $r=2$. We generate 5000 random datasets each consisting of 50 input-output pairs $(\mathbf{x},\mathbf{y})$ where $\mathbf{x}=(x_{1},x_{2})\in\mathbb{R}^{2}$ and $\mathbf{y}=(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}$. For each data pair, $x_{1}$ and $x_{2}$ are sampled uniformly from $[-1,1]$. For $\gamma=1,2,3$, each coordinate $y_{\gamma}$ is a homogeneous polynomial of degree two in two variables $x_{1}$, $x_{2}$, i.e.,
$y_{\gamma}=c_{\gamma,1}x_{1}^{2}+c_{\gamma,2}x_{1}x_{2}+c_{\gamma,3}x_{2}^{2}.$
@@
Figure 3. Comparison of the quadratic polynomials learned by a PNN with $\mathbf{d}=(2,2,3)$ and $r=2$ in each coordinate. Blue points are generated based on the ground truth, while red, green, and yellow stand for the top three most frequent functions learned by the architecture.
For $\gamma=1,2,3$ and $j=1,2,3$, we draw $c_{\gamma,j}$ from a standard normal distribution. We initialize the parameters randomly for each dataset and run gradient descent optimization of the mean squared error (MSE) loss with a learning rate of $0.01$, and decaying by $0.9$ for every $2,000$ training epochs. We stop training at a maximum of ${20,000}$ epochs, and consider the neural networks whose gradient norm satisfies $\|\nabla\ell\|^{2}\leq 10^{-5}$ at the end of training. Instances whose gradient norm does not fall below the threshold within ${20,000}$ iterations are discarded. The final functions learned by the neural networks satisfying the above criteria are approximately critical points. We recover the quadratic coefficients of the function learned by the neural network on each dataset from its weights at epoch ${20,000}$ as follows:
For each output unit $j=1,2,3$,
$c_{1j}^{\prime}$ $=w_{2j1}w_{111}^{2}+w_{2j2}w_{121}^{2},$
$c_{2j}^{\prime}$ $=2(w_{2j1}w_{111}w_{112}+w_{2j2}w_{121}w_{122}),$
$c_{3j}^{\prime}$ $=w_{2j1}w_{112}^{2}+w_{2j2}w_{122}^{2}.$
To assess the distinct functions learned by the neural network, we set $\texttt{tol}=0.1$, and consider two solutions are distinct if there exists $\gamma=1,2,3$ and $j=1,2,3$ such that
$|c_{\gamma j}^{\prime}-\overline{c_{\gamma j}^{\prime}}|\geq\texttt{tol}$
where $c_{\gamma j}^{\prime}$ and $\overline{c_{\gamma j}^{\prime}}$ are corresponding coefficients of the function learned by the neural network by the datasets.
For the architecture $\mathbf{d}=(2,2,3)$ and $r=2$, out of the $5,000$ datasets, $2713$ reached the gradient norm threshold within $20,000$ training epochs. Based on the criteria for the coefficients of the final output, we learned 22 distinct functions whose coefficients differ by the tolerance threshold. We sort the distinct functions by descending the number of training nets it corresponds to. For the top three functions (Function $1$, $2$ and $3$) learned by the architecture, they correspond to $1,557$, $560$, and $241$ datasets, respectively. In the figure below, we evaluate Function $1$ (green), Function $2$ (red), and Function $3$ (yellow) against the ground truth (blue) on the three coordinates $y_{1}$, $y_{2}$, $y_{3}$ separately.
The experiment shows that the three most frequently learned functions all align with the ground truth on the $y_{1}$ and $y_{2}$ coordinates, and being slightly off on $y_{3}$. Note that following the formula in 6.4, the learning degree of the architecture $\mathbf{d}=(2,2,3)$ and $r=2$ is $39$, which provides an upper bound on the number of distinct functions the gradient descent algorithm is able to learn under proper conditions. The experiment implementation is available at [KLW24].
@@ 6.4. Backpropagation
[…]
@@
This routine allows us to compute Jacobians and in particular the dimension of the neurovariety exponentially faster than via the naïve method of computing all derivatives of the parametrization $\Psi_{\mathbf{d},r}$. It can be easily adapted for computations over finite fields which makes dimension computations even faster. In [KTB19] this routine has already been implemented, however, their implementation contains a mistake in the set-up of the linear system described above. We provide a corrected implementation at [KLW24].

### v2

@@ 6.2. Gradient-based optimization
6.3. Experiments
We generated synthetic data consisting of $m=5000$ datasets, each containing $N=50$ data points. For each dataset, input pairs $\hat{\mathbf{x}}=(\hat{x}_{1},\hat{x}_{2})^{\top}$ were sampled uniformly from the range $[-1,1]$. The corresponding outputs $\hat{\mathbf{y}}=(\hat{y}_{1},\hat{y}_{2},\hat{y}_{3})^{\top}$ were generated using quadratic functions with randomly sampled coefficients:
$\hat{y}_{1}$ $=c_{1,1}\hat{x}_{1}^{2}+c_{1,2}\hat{x}_{1}\hat{x}_{2}+c_{1,3}\hat{x}_{2}^{2},$
$\hat{y}_{2}$ $=c_{2,1}\hat{x}_{1}^{2}+c_{2,2}\hat{x}_{1}\hat{x}_{2}+c_{2,3}\hat{x}_{2}^{2},$ (6.7)
$\hat{y}_{3}$ $=c_{3,1}\hat{x}_{1}^{2}+c_{3,2}\hat{x}_{1}\hat{x}_{2}+c_{3,3}\hat{x}_{2}^{2},$
where the coefficients $c_{i,j}$ were sampled from a normal distribution $\mathcal{N}(0,1)$.
We employ a polynomial neural network $p_{\mathbf{w}}$ with architecture $\mathbf{d}=(2,2,3)$ and quadratic activation function $(r=2)$. The parameters $\mathbf{w}$ consist of two weight matrices $\mathbf{w}=(W_{1},W_{2})$ with $W_{1}\in\mathbb{R}^{2\times 2}$ and $W_{2}\in\mathbb{R}^{3\times 2}$.
Each network was trained on its corresponding dataset using stochastic gradient descent (SGD) with an initial learning rate of $\eta=0.1$. The mean squared error (MSE) loss function was minimized:
$L(W_{1},W_{2})=\frac{1}{N}\sum_{i=1}^{N}\lVert p_{\mathbf{w}}(\hat{\mathbf{x}}^{(i)})-\hat{\mathbf{y}}^{(i)}\rVert_{2}^{2}.$
A learning rate scheduler reduced the learning rate by half every $1000$ epochs. Training proceeded for a maximum of $15000$ epochs or until the maximum gradient magnitude fell below a threshold of $1\times 10^{-4}$, indicating convergence. Later, we refer to this hyperparameter as the gradient norm threshold.
After training, we extracted the quadratic coefficients learned by each network to represent the functions it had approximated. The extraction was performed by analytically expressing the network’s output as a quadratic function of the inputs. For each output unit $j$, the coefficients $(a_{1j},a_{2j},a_{3j})$ were computed as
$a_{1j}$ $=v_{j1}w_{11}^{2}+v_{j2}w_{21}^{2},$
$a_{2j}$ $=2(v_{j1}w_{11}w_{12}+v_{j2}w_{21}w_{22}),$ (6.8)
$a_{3j}$ $=v_{j1}w_{12}^{2}+v_{j2}w_{22}^{2},$
where $v_{jk}$ are entries of $W_{2}$ and $w_{kl}$ are entries of $W_{1}$.
To identify distinct functions learned by the network, we compared the extracted coefficients using a coefficient comparison method with a specified tolerance $\epsilon$. Two functions given by coefficients $a_{ij}^{(1)}$ and $a_{ij}^{(2)}$ as in (6.8) were considered the same if
$\left|a_{ij}^{(1)}-a_{ij}^{(2)}\right|<\epsilon\quad\forall i,j.$
We systematically adjusted the tolerance $\epsilon$ to explore its impact on the number of distinct functions identified. By varying $\epsilon$, we could control the granularity of function distinction, allowing minor variations in coefficients to be considered functionally equivalent.
In addition, we performed perturbations on the function’s coefficients to assess whether each learned function corresponded to a local minimum in the neuromanifold. Small perturbations $\delta$ were added to the coefficients,
$\tilde{c}_{ij}=c_{ij}+\delta_{ij},\quad\delta_{ij}\in[-\varepsilon,\varepsilon],$
where $\varepsilon$ is a small value (e.g., $1\times 10^{-4}$). The perturbed coefficients also give rise to a perturbed neural network which we denote by $\tilde{p}_{\mathbf{w}}$. The perturbed loss function then becomes
$L_{\delta}(W_{1},W_{2})=\frac{1}{N}\sum_{i=1}^{N}\left\|\tilde{p}_{\mathbf{w}}(\hat{\mathbf{x}}^{(i)})-\tilde{\mathbf{y}}^{(i)}\right\|_{2}^{2},$
where $\tilde{\mathbf{y}}^{(i)}$ are defined as in (6.7) but using the perturbed coefficients $\tilde{c}_{ij}$. If the value of the original loss function was less than or equal to the losses of all perturbed functions, we considered it to be a local minimum in the neuromanifold.
Through our experimental setup, we were able to
  • train $5000$ datasets with gradient norm with $15000$ epochs and gradient norm threshold $10^{-4}$;
  • identify seventeen distinct functions whose coefficients differ by tolerance threshold $\epsilon=1/10$ and that were learned with frequency no less than $10$ out of $5,000$ datasets;
  • check that for sixteen out of seventeen functions, the coefficient matrices of the functions have rank one;
  • verify the local minimality of the one learned function whose coefficient matrix has rank two in the neuromanifold by evaluating the loss landscape around the function.
The functions with rank-one coefficient matrices correspond to the singular points of the neuromanifold. Since the coefficients of quadratic functions in (6.7) were generated randomly from the normal distribution, we expect the resulting function to be a nonsingular point of the neuromanifold. Therefore we consider only the learned function that is the nonsingular point of the neuromanifold. It is also the most frequently learned function by the network and it is depicted in Figure 3. Having one nonsingular local minimum is consistent with Remark 6.9 that the learning degree is at most three.
@@ 6.3. Experiments
Figure 3. Comparison of the most frequent quadratic polynomial learned by a PNN with $\mathbf{d}=(2,2,3)$ and $r=2$ in each coordinate with the ground truth. The green manifold is generated based on the ground truth, while the red manifold is the most frequent function learned by the network.
The experiment implementation is available at [KLW24].
Remark 6.10.
The initialization of the network parameters poses challenges due to their tendency to suffer from vanishing and exploding gradients. As the depth of a polynomial neural network scales up, the degree of the polynomial increases and the network becomes sensitive to weight variations, causing gradients to either diminish or escalate exponentially during backpropagation. Poor initialization can lead to prolonged training times, convergence to suboptimal solutions, or even complete failure of the training process. While the general optimal initialization scheme remains an open research problem, in these experiments, we carefully initialize the weight and apply gradient clipping to regularize the training process.
@@ 6.4. Backpropagation
Let $a^{l}_{j}$ be the output of the $j^{\text{th}}$ neuron in the $l^{\text{th}}$ layer, and let $L$ denote the last layer of the network. Then the following equations hold:
@@
This routine allows us to compute Jacobians and in particular the dimension of the neurovariety exponentially faster than via the naïve method of computing all derivatives of the parametrization $\Psi_{\mathbf{d},r}$. It can be easily adapted for computations over finite fields which makes dimension computations even faster. In [KTB19] this routine has already been implemented, however (as of Oct 9 2024), their implementation at [KTB] contains a mistake in the set-up of the linear system described above. We provide a corrected implementation at [KLW24].
@@ 6.5. Case study: Linear Neural Networks
7. Conclusion
We have studied the functional space of neural networks with polynomial activation functions from an algebro-geometric perspective. These networks, characterized by their ability to model complex, high-degree polynomial relationships in data, provide a unique lens through which the intricate dynamics of deep learning models can be better understood. Moreover, since a large class of functions can be efficiently approximated by polynomials, delving into the polynomial framework can be important for the development of more efficient and accurate models.
In this paper, we have focused on the geometry of polynomial neural networks. We have characterized neuromanifolds and neurovarieties for some specific architectures, and studied their dimensions. The geometry and dimension of the neuromanifold relates directly to the expressive power of the neuromanifold. Beyond expressivity, we also studied the optimization process of polynomial neural networks, and bound the number of critical points of a loss function using the learning degree. Besides our conjectures from §5 concerning the dimension of neurovarieties, there are further questions to be studied, for example: What are the learning degrees of more general architectures? How many of the critical points of the loss functions are real?